# Solvability of fractional semilinear heat equations with distributional forcing terms in Lorentz spaces

### v1

@@ 2.1 Basic properties of Lorentz spaces
$\alpha_{(f\ast g)\chi_{B(z,1)}}(\lambda)$ $\leqslant\alpha_{(f_{1}\ast g)\chi_{B(z,1)}}\left(\frac{\lambda}{2}\right)\leqslant\frac{2}{\lambda}\left\|f_{1}\ast g\mathrel{}\middle|\mathrel{}L^{1}(B(z,1))\right\|$
$\leqslant\frac{2}{\lambda}\int_{B(z,1)}\int_{\mathbb{R}^{N}}\left|f_{1}(x-y)\right||g(y)|\,dydx$
$=\frac{2}{\lambda}\int_{\mathbb{R}^{N}}|g(y)|\int_{B(z+y,1)}\left|f_{1}(x)\right|\,dxdy.$ (2.1)
@@
$\int_{B(z+y,1)}\left|f_{1}(x)\right|\,dx$ $=\int_{B(z+y,1)}\int_{0}^{\infty}\chi_{\left\{0<\lambda<|f_{1}(x)|\right\}}(\lambda)\,d\lambda dx=\int_{0}^{\infty}\alpha_{f_{1}\chi_{B(z+y,1)}}(\lambda)\,d\lambda$
$=\int_{0}^{M}\alpha_{f\chi_{B(z+y,1)}}(M)\,d\lambda+\int_{M}^{\infty}\alpha_{f\chi_{B(z+y,1)}}(\lambda)\lambda^{p}\lambda^{-p}\,d\lambda$
$\leqslant M^{1-p}\left\|f\chi_{B(z+y,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|^{p}+\int_{M}^{\infty}\frac{d\lambda}{\lambda^{p}}\left\|f\chi_{B(z+y,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|^{p}$
$\leqslant\frac{p}{p-1}M^{1-p}\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{p}.$ (2.2)
We plug (2.1) into (2.1) and recall $M=\frac{\lambda}{2\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|}$ to see that
$\lambda\alpha_{(f\ast g)\chi_{B(z,1)}}(\lambda)^{1/p}$ $\leqslant\lambda\left[\frac{2}{\lambda}\int_{\mathbb{R}^{N}}|g(y)|\frac{p}{p-1}\left(\frac{\lambda}{2\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|}\right)^{1-p}\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{p}\,dy\right]^{1/p}$
$\leqslant 2\left(\frac{p}{p-1}\right)^{1/p}\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
holds for all $\lambda>0$ and $z\in\mathbb{R}^{N}$. Taking the sup in $\lambda$ and $z$ yields the claim. ∎
2.2 Real interpolation of Lorentz spaces
@@
$\left(X_{0},X_{1}\right)_{\kappa,q}:=\left\{f\in X_{0}+X_{1}\mathrel{}\middle|\mathrel{}\left\|f\mathrel{}\middle|\mathrel{}\left(X_{0},X_{1}\right)_{\kappa,q}\right\|<\infty\right\}$
, where
@@
Let $p_{0}$ and $p_{1}\in\left]0,\infty\right[$ satisfy $p_{0}\neq p_{1}$ and take $q\in\left]0,\infty\right]$ and $\kappa\in\left]0,1\right[$. We set $p\in\left]0,\infty\right[$ by the formula $p^{-1}=(1-\kappa)p_{0}^{-1}+\kappa p_{1}^{-1}$. Then, we have $\left(L^{p_{0},q_{0}},L^{p_{1},q_{1}}\right)_{\kappa,q}=L^{p,q}$ for all $q_{0},q_{1}\in\left]0,\infty\right]$.
@@
with $p_{0},p_{1}\in\left]1,\infty\right[$ by means of Theorem 2.6. Thus, Proposition 2.5 (1) and (2) reveal the space $L^{p,q}$ as a Banach space.
It is useful for Section 3 to prepare a $L^{p,q}_{\mathrm{ul}}$ version of Theorem 2.6.
Corollary 2.8.
Let $p_{0}$, $p_{1}$, $q$, $\kappa$ and $p$ be as in Theorem 2.6. Then, the embedding
$\left(L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}}\right)_{\kappa,q}\subset L^{p,q}_{\mathrm{ul}}$
is valid for all $q_{0}$, $q_{1}\in\left]0,\infty\right]$.
Proof.
Fix $z\in\mathbb{R}^{N}$ and $\lambda>0$. Take $f\in\left(L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}}\right)_{\kappa,q}$. Define $A$ and $B(\subset\mathbb{R})$ as
$A(z,\lambda,f)$ $:=\left\{\left\|f_{0}\mathrel{}\middle|\mathrel{}L^{p_{0},q_{0}}\right\|+\lambda\left\|f_{1}\mathrel{}\middle|\mathrel{}L^{p_{1},q_{1}}\right\|\mathrel{}\middle|\mathrel{}\begin{subarray}{c}f\chi_{B(z,1)}=f_{0}+f_{1},\\ f_{0}\in L^{p_{0},q_{0}},\,\,f_{1}\in L^{p_{1},q_{1}}\end{subarray}\right\},$
$B(z,\lambda,f)$ $:=\left\{\left\|f_{0}\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p_{0},q_{0}}\right\|+\lambda\left\|f_{1}\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p_{1},q_{1}}\right\|\mathrel{}\middle|\mathrel{}\begin{subarray}{c}f=f_{0}+f_{1},\\ f_{0}\in L^{p_{0},q_{0}}_{\mathrm{ul}},\qquad f_{1}\in L^{p_{1},q_{1}}_{\mathrm{ul}}\end{subarray}\right\}.$
We can easily check $A(z,\lambda,f)\supset B(z,\lambda,f)$ and this inclusion implies that
$K(f\chi_{B(z,1)},\lambda;L^{p_{0},q_{0}},L^{p_{1},q_{1}})=\inf A(z,\lambda,f)\leqslant\inf B(z,\lambda,f)$
$\leqslant\inf\left\{\left\|f_{0}\mathrel{}\middle|\mathrel{}L^{p_{0},q_{0}}_{\mathrm{ul}}\right\|+\lambda\left\|f_{1}\mathrel{}\middle|\mathrel{}L^{p_{1},q_{1}}_{\mathrm{ul}}\right\|\mathrel{}\middle|\mathrel{}\begin{subarray}{c}f=f_{0}+f_{1},\\ f_{0}\in L^{p_{0},q_{0}}_{\mathrm{ul}},\qquad f_{1}\in L^{p_{1},q_{1}}_{\mathrm{ul}}\end{subarray}\right\}=K(f,\lambda;L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}})$
for all $z\in\mathbb{R}^{N}$ and $\lambda>0$. In the light of Theorem 2.6, we have
@@
$\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p,q}\right\|$ $\leqslant C\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}\left(L^{p_{0},q_{0}},L^{p_{1},q_{1}}\right)_{\kappa,q}\right\|$
@@
$\leqslant C\begin{dcases}\left(\int_{0}^{\infty}\left(\lambda^{-\kappa}K(f,\lambda;L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}})\right)^{q}\,\frac{d\lambda}{\lambda}\right)^{1/q}&\text{for }q\in\left]0,\infty\right[,\\ \sup_{\lambda>0}\,\lambda^{-\kappa}K(f,\lambda;L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}})&\text{for }q=\infty\end{dcases}$
@@
$=C\left\|f\mathrel{}\middle|\mathrel{}\left(L^{p_{0},q_{0}}_{\mathrm{ul}},L^{p_{1},q_{1}}_{\mathrm{ul}}\right)_{\kappa,q}\right\|.$
@@
Taking the supremum over $z\in\mathbb{R}^{N}$ leads the desired embedding. ∎
2.3 Properties of Besov–Lorentz spaces
We prove some properties of Besov–Lorentz spaces, with reference to [17].
@@ 2.3 Properties of Besov–Lorentz spaces
$B^{0}_{(p,\infty),1}\subset L^{p,\infty}_{\mathrm{ul}}\subset B^{0}_{(p,\infty),\infty}$
for $p\in\left]1,\infty\right[$.
@@
By using Proposition 2.3, we have
$\left\|\Delta_{j}f\mid L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant\begin{cases}C\left\|\mathcal{F}^{-1}\varphi_{j}\mid L^{1}\right\|\left\|f\mid L^{p,\infty}_{\mathrm{ul}}\right\|=C\left\|\mathcal{F}^{-1}\varphi_{0}\mid L^{1}\right\|\left\|f\mid L^{p,\infty}_{\mathrm{ul}}\right\|&\text{for }j\geqslant 1\\ C\left\|\mathcal{F}^{-1}\varphi_{(0)}\mid L^{1}\right\|\left\|f\mid L^{p,\infty}_{\mathrm{ul}}\right\|&\text{for }j=0\end{cases}$
for all $f\in L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})$. Taking the supremum over $j\in\mathbb{N}$ gives the second inclusion. Next we take $f\in B^{0}_{(p,\infty),1}$. We can easily check that $f=\sum_{j\in\mathbb{N}}\Delta_{j}f$ in $\mathcal{S}^{\prime}(\mathbb{R}^{N})$ and can conduct the following calculation
$\left\|\sum_{j=0}^{n}\Delta_{j}f-\sum_{j=0}^{m}\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant\sum_{j=m+1}^{n}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\left(\leqslant\left\|f\mid B^{0}_{(p,\infty),1}\right\|<\infty\right)$
@@
Thus $\sum_{j\in\mathbb{N}}\Delta_{j}f=f$ belongs to $L^{p,\infty}_{\mathrm{ul}}$. The continuity is easily seen. ∎
Note that we didn’t use properties of Lorentz spaces apart from their completeness and Proposition 2.3 (see also Proposition 2.11 in [17]). We next verify the embedding theorem of Sobolev-type.
@@
(Sobolev-type embedding) Let $1\leqslant r\leqslant\infty$, $s\in\mathbb{R}$, $1<p<\infty$ and $1\leqslant q\leqslant\infty$. We have the following continuous embeddings:
@@
    $B^{s}_{(p,q),r}\subset B^{s-N(1-\theta)/p}_{\left(p/\theta,q/\theta\right),r}$ for $\theta\in\left]0,1\right[$.
Here, $B^{s-N/p}_{P,r}$ stands for inhomogeneous Besov spaces defined in the same way as inhomogeneous Besov–Lorentz spaces by replacing Lorentz spaces with standard $L^{P}$ space.
To obtain this, we modify the proof of Theorem 2.5 in [17]. We mention that there is another way to prove this claim (see Remark 2.11 below).
Proof of Proposition 2.10.
First we show the following inequality
$\sup_{x\in\mathbb{R}^{N}}\left|\Delta_{j}f(x)\right|\leqslant C2^{jN/p}\left\|\Delta_{j}f\mid L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant C2^{jN/p}\left\|\Delta_{j}f\mid L^{p,q}_{\mathrm{ul}}\right\|$ (2.3)
is valid for all $f\in L^{p,q}_{\mathrm{ul}}$ and for all $j\in\mathbb{N}$ (the latter is a standard fact).
We write $\Phi_{j}(\xi):=\varphi_{j-1}(\xi)+\varphi_{j}(\xi)+\varphi_{j+1}(\xi)$ for $j\in\mathbb{Z}$ and $\Phi_{(0)}(\xi):=\varphi_{(0)}(\xi)+\varphi_{1}(\xi)$. Then identities $\varphi_{j}(\xi)\equiv\Phi_{j}(\xi)\varphi_{j}(\xi)$, $\varphi_{(0)}(\xi)\equiv\Phi_{(0)}(\xi)\varphi_{(0)}(\xi)$ and $\Psi_{j}(x):=\mathcal{F}^{-1}\Phi_{j}(x)=2^{Nj}\mathcal{F}^{-1}\Phi_{0}(2^{j}x)$ hold. Fix $x\in\mathbb{R}^{N}$. We decompose $|\Delta_{j}f(x)|$ as
@@
$\leqslant 2^{Nj}\left(\int_{B(0,1)}+\sum_{k=0}^{\infty}\int_{B(0,2^{k+1})\setminus B(0,2^{k})}\right)\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(x-y)\right|\,dy.$ (2.4)
@@
$2^{Nj}\int_{B(0,1)}\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(x-y)\right|\,dy\leqslant 2^{Nj}\int_{\mathbb{R}^{N}}\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(y)\chi_{B(x,1)}(y)\right|\,dy$
$\leqslant$  $\,2^{Nj}\left\|\Psi_{0}(2^{j}y)\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\chi_{B(x,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|\leqslant 2^{Nj/p}\left\|\Psi_{0}\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$.
Note that there exists a positive integer $\Omega=\Omega_{N}$ such that
$B(0,2)\subset\bigcup_{m=1}^{\Omega}B(z_{m},1)\quad\text{for some}\quad\left\{z_{m}\right\}_{m=1}^{\Omega}$
and therefore we can pick $\left\{z_{m}\right\}_{m=1}^{\Omega^{k}}$ satisfying $B(0,2^{k})\subset\bigcup_{m=1}^{\Omega^{k}}B(z_{m},1)$ for each $k\in\mathbb{N}$. Since $\Psi_{0}=\mathcal{F}^{-1}\Phi_{0}$ is a rapidly decreasing function, it holds that

### v2

@@ 2.1 Basic properties of Lorentz–Morrey spaces
$\alpha_{(f\ast g)\chi_{B(z,R)}}(\lambda)$ $\leqslant\alpha_{(f_{1}\ast g)\chi_{B(z,R)}}\left(\frac{\lambda}{2}\right)\leqslant\frac{2}{\lambda}\left\|f_{1}\ast g\mathrel{}\middle|\mathrel{}L^{1}(B(z,R))\right\|$
$\leqslant\frac{2}{\lambda}\int_{B(z,R)}\int_{\mathbb{R}^{N}}\left|f_{1}(x-y)\right||g(y)|\,dy\,dx$
$=\frac{2}{\lambda}\int_{\mathbb{R}^{N}}|g(y)|\int_{B(z-y,R)}\left|f_{1}(x)\right|\,dx\,dy.$ (2.1)
@@
$\int_{B(z-y,R)}\left|f_{1}(x)\right|\,dx$ $=\int_{B(z-y,R)}\int_{0}^{\infty}\chi_{\left\{0<k<|f_{1}(x)|\right\}}(k)\,dk\,dx=\int_{0}^{\infty}\alpha_{f_{1}\chi_{B(z-y,R)}}(k)\,dk$
$=\int_{0}^{M}\alpha_{f\chi_{B(z-y,R)}}(M)\,dk+\int_{M}^{\infty}\alpha_{f\chi_{B(z-y,R)}}(k)k^{q}k^{-q}\,dk$
$\leqslant M^{1-q}\left\|f\chi_{B(z-y,R)}\mathrel{}\middle|\mathrel{}L^{q,\infty}\right\|^{q}+\int_{M}^{\infty}\frac{dk}{k^{q}}\left\|f\chi_{B(z-y,R)}\mathrel{}\middle|\mathrel{}L^{q,\infty}\right\|^{q}$
$\leqslant\frac{q}{q-1}M^{1-q}R^{N-\frac{Nq}{p}}\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|^{q}.$ (2.2)
We plug (2.1) into (2.1) and recall $M=\lambda/2\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|$ to see that
$R^{\frac{N}{p}-\frac{N}{q}}\lambda\alpha_{(f\ast g)\chi_{B(z,R)}}(\lambda)^{1/q}$ $\leqslant 2\left(\frac{q}{q-1}\right)^{1/q}\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
holds for all $R\in\left]0,1\right]$, $z\in\mathbb{R}^{N}$ and $\lambda>0$. Taking the sup in $\lambda>0$ yields
$R^{\frac{N}{p}-\frac{N}{q}}\left\|(f\ast g)\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{q,\infty}\right\|$ $\leqslant 2\left(\frac{q}{q-1}\right)^{1/q}\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
for all $(z,R)\in\mathbb{R}^{N}\times\left]0,1\right]$, and thus we arrive at the claim. ∎
2.2 Real interpolation of Lorentz–Morrey spaces
@@
$\left(X_{0},X_{1}\right)_{\kappa,q}:=\left\{f\in X_{0}+X_{1}\mathrel{}\middle|\mathrel{}\left\|f\mathrel{}\middle|\mathrel{}\left(X_{0},X_{1}\right)_{\kappa,q}\right\|<\infty\right\},$
where
@@
Let $p_{0}$ and $p_{1}\in\left]0,\infty\right[$ satisfy $p_{0}\neq p_{1}$ and take $q\in\left]0,\infty\right]$ and $\kappa\in\left]0,1\right[$. We set $p\in\left]0,\infty\right[$ by the formula $p^{-1}=(1-\kappa)p_{0}^{-1}+\kappa p_{1}^{-1}$. Then we have $\left(L^{p_{0},q_{0}},L^{p_{1},q_{1}}\right)_{\kappa,q}=L^{p,q}$ for all $q_{0},q_{1}\in\left]0,\infty\right]$.
@@
with $p_{0},p_{1}\in\left]1,\infty\right[$, by means of Theorem 2.6. Thus, Proposition 2.5 (1) and (2) identify the space $L^{p,q}$ as a Banach space.
We prepare a real interpolation result for Lorentz–Morrey spaces.
Proposition 2.8.
@@
Let $\left(p_{0},p_{1},q_{0},q_{1}\right)\in\left]0,\infty\right[^{4}$ satisfy $p_{0}\neq p_{1}$, $q_{0}\neq q_{1}$, $q_{0}\leqslant p_{0}$ and $q_{1}\leqslant p_{1}$. We take $r\in\left]0,\infty\right]$ and $\kappa\in\left]0,1\right[$, and set $p$, $q$ by the formula
$\left(p^{-1},q^{-1}\right)=(1-\kappa)\left(p_{0}^{-1},q_{0}^{-1}\right)+\kappa\left(p_{1}^{-1},q_{1}^{-1}\right).$
Then $q\leqslant p$ holds and the embedding
$\left(M^{p_{0}}_{q_{0},r_{0}},M^{p_{1}}_{q_{1},r_{1}}\right)_{\kappa,r}\subset M^{p}_{q,r}$
is valid for all $r_{0}$, $r_{1}\in\left]0,\infty\right]$.
This property has a key role in section 3. It is noteworthy that this kind of embedding fails in classical Morrey spaces (see Theorem 3 (v) in [23]). We prove this property in a more general form, for which we introduce the following definition.
Definition 2.9.
($X$-Morrey spaces) Let $X$ be a quasi-normed function space and $p\in\left]0,\infty\right[$. The local $X$-Morrey space $M^{p}_{X}$ is defined as the set of functions $f\in X$ such that $\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{X}\right\|<\infty$, where
$\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{X}\right\|:=\sup_{0<R\leqslant 1}\sup_{z\in\mathbb{R}^{N}}R^{\frac{N}{p}}\left\|\left(f_{R}\right)\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}X\right\|.$
Here, we set $f_{\lambda}(x):=f(\lambda x)$ for all $\lambda>0$.
Lemma 2.1.(4) suggests that this definition is consistent with our Lorentz–Morrey spaces. Note that $\left(f_{R}\right)\chi_{B(z,1)}=\left(f\chi_{B(Rz,R)}\right)_{R}$ holds for all $R>0$ and $z\in\mathbb{R}^{N}$.
Proposition 2.10.
Let $(p_{0},p_{1})\in\left]0,\infty\right[^{2}$ satisfy $p_{0}\neq p_{1}$. We take $r\in\left]0,\infty\right]$ and $\kappa\in\left]0,1\right[$, and set $p$ by the formula $p^{-1}=(1-\kappa)p_{0}^{-1}+\kappa p_{1}^{-1}$. Assume that quasi-normed function spaces $X$, $X_{0}$, $X_{1}$ have the property $\left(X_{0},X_{1}\right)_{\kappa,r}\subset X.$ Then the embedding $\left(M^{p_{0}}_{X_{0}},M^{p_{1}}_{X_{1}}\right)_{\kappa,r}\subset M^{p}_{X}$ is valid.
Proof of Proposition 2.10.
We restrict ourselves to the case $r=\infty$ (the remaining cases are analogous). Fix $z\in\mathbb{R}^{N}$ and $R\in\left]0,1\right]$. Take $f\in\left(M^{p_{0}}_{X_{0}},M^{p_{1}}_{X_{1}}\right)_{\kappa,\infty}$. Let $(g,h)\in M^{p_{0}}_{X_{0}}\times M^{p_{1}}_{X_{1}}$ be an arbitrary decomposition $f=g+h$. Then the pair $\left(g_{R}\chi_{B(z,1)},h_{R}\chi_{B(z,1)}\right)$ becomes an element of $X_{0}\times X_{1}$ and satisfies $f_{R}\chi_{B(z,1)}=g_{R}\chi_{B(z,1)}+h_{R}\chi_{B(z,1)}$. Thus, for all fixed $\lambda>0$, we have
$K\left(f_{R}\chi_{B(z,1)},\lambda;X_{0},X_{1}\right)$ $\leqslant\left\|g_{R}\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}X_{0}\right\|+\lambda\left\|h_{R}\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}X_{1}\right\|$
$\leqslant R^{-\frac{N}{p_{0}}}\left\|g\mathrel{}\middle|\mathrel{}M^{p_{0}}_{X_{0}}\right\|+\lambda R^{-\frac{N}{p_{1}}}\left\|h\mathrel{}\middle|\mathrel{}M^{p_{1}}_{X_{1}}\right\|$
$=R^{-\frac{N}{p_{0}}}\left\{\left\|g\mathrel{}\middle|\mathrel{}M^{p_{0}}_{X_{0}}\right\|+\lambda R^{\frac{N}{p_{0}}-\frac{N}{p_{1}}}\left\|h\mathrel{}\middle|\mathrel{}M^{p_{1}}_{X_{1}}\right\|\right\}.$
Setting $\Lambda:=\lambda R^{\frac{N}{p_{0}}-\frac{N}{p_{1}}}$, the arbitrariness of the decomposition implies
$K\left(f_{R}\chi_{B(z,1)},\lambda;X_{0},X_{1}\right)$ $\leqslant R^{-\frac{N}{p_{0}}}\Lambda^{\kappa}\Lambda^{-\kappa}K\left(f,\Lambda;M^{p_{0}}_{X_{0}},M^{p_{1}}_{X_{1}}\right)$
@@
$\leqslant R^{-\frac{N}{p_{0}}}\Lambda^{\kappa}\left\|f\mathrel{}\middle|\mathrel{}\left(M^{p_{0}}_{X_{0}},M^{p_{1}}_{X_{1}}\right)_{\kappa,\infty}\right\|$
for all $\lambda>0$. Therefore, in the light of $X\supset\left(X_{0},X_{1}\right)_{\kappa,\infty}$, we obtain
@@
$R^{\frac{N}{p}}\left\|f_{R}\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}X\right\|$ $\leqslant CR^{\frac{N}{p}}\left\|f_{R}\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}\left(X_{0},X_{1}\right)_{\kappa,\infty}\right\|$
$=CR^{\frac{N}{p}}\sup_{\lambda>0}\lambda^{-\kappa}K\left(f_{R}\chi_{B(z,1)},\lambda;X_{0},X_{1}\right)$
$\leqslant CR^{\frac{N}{p}}\sup_{\lambda>0}\lambda^{-\kappa}R^{-\frac{N}{p_{0}}}\Lambda^{\kappa}\left\|f\mathrel{}\middle|\mathrel{}\left(M^{p_{0}}_{X_{0}},M^{p_{1}}_{X_{1}}\right)_{\kappa,\infty}\right\|$
$=C\left\|f\mathrel{}\middle|\mathrel{}\left(M^{p_{0}}_{X_{0}},M^{p_{1}}_{X_{1}}\right)_{\kappa,\infty}\right\|.$
@@
Taking the supremum over $z\in\mathbb{R}^{N}$ and $R\in\left]0,1\right]$ leads the desired embedding. ∎
Proof of Proposition 2.8.
Clearly, $q\leqslant p$. Use Theorem 2.6 and Proposition 2.10 with $X_{i}=L^{q_{i},r_{i}}$ ($i\in\left\{,0,1\right\}$). ∎
2.3 Properties of BLM spaces
We prove some properties of BLM spaces, with reference to [20].
@@ 2.3 Properties of BLM spaces
$B^{0,1}_{p,q,\infty}\subset M^{p}_{q,\infty}\subset B^{0,\infty}_{p,q,\infty}$
for $1<q\leqslant p<\infty$.
@@
Proposition 2.3 and $\left\|\mathcal{F}^{-1}\varphi_{j}\mathrel{}\middle|\mathrel{}L^{1}\right\|=\left\|\mathcal{F}^{-1}\varphi_{0}\mathrel{}\middle|\mathrel{}L^{1}\right\|$ (for $j\in\mathbb{Z}$) give that
$\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant\begin{cases}C\left\|\mathcal{F}^{-1}\varphi_{0}\mid L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|&\text{for }j\geqslant 1\\ C\left\|\mathcal{F}^{-1}\varphi_{(0)}\mid L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|&\text{for }j=0\end{cases}$
holds for all $f\in M^{p}_{q,\infty}(\mathbb{R}^{N})$. We take the supremum over $j\in\mathbb{N}$ to get the second one. Next we fix $f\in B^{0,1}_{p,q,\infty}$. We can easily check that $f=\sum_{j\in\mathbb{N}}\Delta_{j}f$ in $\mathcal{S}^{\prime}(\mathbb{R}^{N})$ and can conduct the following calculation
$\left\|\sum_{j=0}^{n}\Delta_{j}f-\sum_{j=0}^{m}\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ $\leqslant\sum_{j=m+1}^{n}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\left(\leqslant\left\|f\mid B^{0,1}_{p,q,\infty}\right\|<\infty\right)$
@@
It turns out that $f\in M^{p}_{q,\infty}$. The embedding is easily seen to be continuous. ∎
Note that we didn’t use properties of Lorentz–Morrey spaces apart from their completeness and Proposition 2.3 (see also Proposition 2.11 in [20]). We next verify the embedding theorem of Sobolev-type.
@@
(Sobolev-type embedding) Let $r,\kappa\in\left[1,\infty\right]$, $1<q\leqslant p<\infty$ and $s\in\mathbb{R}$. We have the following continuous embeddings:
@@
    $B^{s,r}_{p,q,\kappa}\subset B^{s-N(1-l)/p,r}_{p/l,q/l,\kappa}$ for $l\in\left]0,1\right[$.
Here, $B^{s-N/p,r}_{P}$ stands for inhomogeneous Besov spaces defined in the same way as inhomogeneous BLM spaces by replacing Lorentz–Morrey spaces with the standard $L^{P}$ space.
To obtain this, we modify the proof of Theorem 2.5 in [20]. First we prepare the following estimate.
Lemma 2.13.
Let $1<q\leqslant p<\infty$ and $1\leqslant\kappa\leqslant\infty$. There exists $C>0$ such that the inequalities
$\sup_{x\in\mathbb{R}^{N}}\left|\Delta_{j}f(x)\right|\leqslant C2^{jN/p}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant C2^{jN/p}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\kappa}\right\|$ (2.3)
are valid for all $f\in M^{p}_{q,\infty}$ and for all $j\in\mathbb{N}$.
We prove only the former inequality (the latter one is a consequence of the standard fact that $L^{q,\kappa}\subset L^{q,\infty}$).
Proof of Lemma 2.13.
We write $\Phi_{j}(\xi):=\varphi_{j-1}(\xi)+\varphi_{j}(\xi)+\varphi_{j+1}(\xi)$ for $j\in\mathbb{Z}$ and $\Phi_{(0)}(\xi):=\varphi_{(0)}(\xi)+\varphi_{1}(\xi)$. Then identities $\varphi_{j}(\xi)\equiv\Phi_{j}(\xi)\varphi_{j}(\xi)$, $\varphi_{(0)}(\xi)\equiv\Phi_{(0)}(\xi)\varphi_{(0)}(\xi)$ and $\Psi_{j}(x):=\mathcal{F}^{-1}\Phi_{j}(x)=2^{Nj}\Psi_{0}(2^{j}x)$ hold. Fix $x\in\mathbb{R}^{N}$. We decompose $|\Delta_{j}f(x)|$ as
@@
$\leqslant 2^{Nj}\left(\int_{B(0,2^{-j})}+\sum_{k=-j}^{\infty}\int_{B(0,2^{k+1})\setminus B(0,2^{k})}\right)\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(x-y)\right|\,dy.$ (2.4)
@@
$\int_{B(0,2^{-j})}\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(x-y)\right|\,dy\leqslant\int_{\mathbb{R}^{N}}\left|\Psi_{0}(2^{j}y)\right|\left|\Delta_{j}f(x-y)\chi_{B(0,2^{-j})}(y)\right|\,dy$
$\leqslant$ $\,\left\|\Psi_{0}(2^{j}y)\mathrel{}\middle|\mathrel{}L^{\frac{q}{q-1},1}\right\|\left\|\Delta_{j}f\chi_{B(x,2^{-j})}\mathrel{}\middle|\mathrel{}L^{q,\infty}\right\|\leqslant C2^{\frac{Nj(1-p)}{p}}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|.$
Note that there exists a positive integer $\Omega=\Omega_{N}$ such that for all $r>0$,
$B(0,2r)\subset\bigcup_{m=1}^{\Omega}B(z_{m},r)\quad\text{holds for some}\quad\left\{z_{m}\right\}_{m=1}^{\Omega},$
and therefore we can pick $\left\{z_{m}\right\}_{m=1}^{\Omega^{k}}$ satisfying $B(0,2^{k}r)\subset\bigcup_{m=1}^{\Omega^{k}}B(z_{m},r)$ for each $k\in\mathbb{N}$ and for all $r>0$. Since $\Psi_{0}=\mathcal{F}^{-1}\Phi_{0}$ is a rapidly decreasing function, it holds that